Circular No. 8243

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003kd AND 2003ke Further to IAUC 8242, B. Swift, R. Foley, and W. Li report the LOSS

discovery of two apparent supernovae on unfiltered KAIT images:

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003kd Nov. 23.3  3^{h}00^{m}45.^{s}80 +44^{o}27′13′′.0 17.1 0′′.6 W, 6′′.7 N
2003ke Nov. 23.6  9 45 52.96 +34 41 01.4 18.2 9′′.3 E, 7′′.0 S

Additional approximate magnitudes: SN 2003kd in UGC 2468, Oct. 26.3 UT, [19.5; Nov. 24.4, 16.7. SN 2003ke in MCG +06-22-9, Apr. 4.2, [19.5; Nov. 25.5, 18.3.

SUPERNOVAE 2003kb AND 2003kc A. V. Filippenko, R. J. Foley, and B. J. Swift write that inspection of

CCD spectra obtained at Lick, as noted on IAUC 8242, on Nov. 23 UT reveals that two of the classifications should be revised; the new spectrograms have considerably higher quality than did those obtained on Nov. 22, when conditions were poor. SN 2003kb is probably of type Ic, ∼ 1 week past maximum brightness (rather than type Ia, ∼ 2 weeks past maximum). SN 2003kc is a relatively normal type-Ia supernova near maximum brightness; its spectrum is not very similar to that of SN 1991T.

COMET C/2003 U1 (LINEAR) Improved elliptical orbital elements from MPEC 2003-W57:

T = 2003 Nov. 3.3973 TT  ω = 278.^{o}3829
e = 0.921616  Ω = 322.7625 }2000.0
q = 1.795719 AU  i = 164.4745

a = 22.909194 AU n^{o} = 0.0089885 P = 110 years

COMET C/2002 T7 (LINEAR) Visual total-magnitude estimates: Sept. 22.71 UT, 11.5 (S. Yoshida,

Ibaraki, Japan, 0.25-m reflector); Oct. 5.05, 11.7 (B. H. Granslo, Fjellhamar, Norway, 0.20-m reflector); 27.95, 10.8 (M. Meyer, Kelkheim, Germany, 20×100 binoculars); Nov. 9.05, 9.8 (T. Scarmato, Calabria, Italy, 0.25-m reflector); 17.88, 10.1 (J. J. Gonzalez, Leon, Spain, 25×100 binoculars); 24.19, 9.5 (C. S. Morris, Fillmore, CA, 20×80 binoculars).

2003 November 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green